OpenStudy (anonymous):

If f(x) = [ 4 tanx / x ], find f′( x ).

OpenStudy (anonymous):

\[f(x)=\frac{4tanx}{x} => f'(x)=\frac{ x(4tanx)'-4tanx(x)' }{ (x)^2 }=\frac{ 4xsec ^2 x-4\tan x }{x^2 }\]